# Att-Adapter: A Robust and Precise Domain-Specific Multi-Attributes T2I Diffusion Adapter via Conditional Variational Autoencoder

### v2

@@ Abstract
Text-to-Image (T2I) Diffusion Models have achieved remarkable performance in generating high quality images. However, enabling precise control of continuous attributes, especially multiple attributes simultaneously, in a new domain (e.g., numeric values like eye openness or car width) with text-only guidance remains a significant challenge. To address this, we introduce the Attribute (Att) Adapter, a novel plug-and-play module designed to enable fine-grained, multi-attributes control in pretrained diffusion models. Our approach learns a single control adapter from a set of sample images that can be unpaired and contain multiple visual attributes. The Att-Adapter leverages the decoupled cross attention module to naturally harmonize the multiple domain attributes with text conditioning. We further introduce Conditional Variational Autoencoder (CVAE) to the Att-Adapter to mitigate overfitting, matching the diverse nature of the visual world. Evaluations on two public datasets show that Att-Adapter outperforms all LoRA-based baselines in controlling continuous attributes. Additionally, our method enables a broader control range and also improves disentanglement across multiple attributes, surpassing StyleGAN-based techniques. Notably, Att-Adapter is flexible, requiring no paired synthetic data for training, and is easily scalable to multiple attributes within a single model.
@@ 4.4 Analysis and Application
Exploring training settings. We explore several different settings of our method. The details are in Sec. C in the appendix. Some of the takeaways from the experiments are as follows. 1) Model performance is converged around 200k iterations with batch size of 4. 2) Embedding dimension d (i.e., $Z\in\mathbb{R}^{\text{d}}$) needs to be bigger than 128 such as 512 or 2048. 3) The performance difference between isotropic and non-isotropic gaussian prior is marginal.
Controlling multiple attributes. Fig. 4 and Fig. 10 shows the superior performance of Att-Adapter in controlling multi-attributes. More examples and a detailed explanation about the additional attributes yaw and pitch are provided in the appendix (Sec. E.1).
@@
We also show the results of combining LoRA finetuned for a person ID with Att-Adapter in the appendix (Sec. B).
@@ Appendix A Evaluation Details
Appendix B Application results: LoRA
@@
Appendix C Exploring training settings.
@@
Appendix D Additional comparison of Att-Adapter and LoRA
@@
Appendix E Additional Analysis
E.1 Challenging attributes
In this section, we show that Att-Adapter can be used beyond frontal-view images. In order to show this, we extend our training dataset from the frontal-view face images to the entire FFHQ dataset. We also additionally add three attributes; yaw, pitch, and roll [^6] https://github.com/DCGM/ffhq-features-dataset. The results are shown in Fig. 15. Interestingly, even though Att-Adapter is not specifically designed for understanding 3D domain, we can see that the left-right rotation (i.e., yaw) and the up-down rotation (i.e., pitch) can be controlled. However, we observed that roll is not controllable. To improve this, we believe additional facial dataset with diverse roll information is required as FFHQ face-cropped dataset is face aligned. We also think it would be interesting and powerful if 3D domain knowledge could be aggregated in Att-Adapter which is beyond our research scope.

### v4

@@ Abstract
Text-to-Image (T2I) Diffusion Models have achieved remarkable performance in generating high quality images. However, enabling precise control of continuous attributes, especially multiple attributes simultaneously, in a new domain (e.g., numeric values like eye openness or car width) with text-only guidance remains a significant challenge. To address this, we introduce the Attribute (Att) Adapter, a novel plug-and-play module designed to enable fine-grained, multi-attributes control in pretrained diffusion models. Our approach learns a single control adapter from a set of sample images that can be unpaired and contain multiple visual attributes. The Att-Adapter leverages the decoupled cross attention module to naturally harmonize the multiple domain attributes with text conditioning. We further introduce Conditional Variational Autoencoder (CVAE) to the Att-Adapter to mitigate overfitting, matching the diverse nature of the visual world. Evaluations on two public datasets show that Att-Adapter outperforms all LoRA-based baselines in controlling continuous attributes. Additionally, our method enables a broader control range and also improves disentanglement across multiple attributes, surpassing StyleGAN-based techniques. Notably, Att-Adapter is flexible, requiring no paired synthetic data for training, and is easily scalable to multiple attributes within a single model. The project page is available at: https://tri-mac.github.io/att-adapter/.
@@ 2 Related Works
Additional related works on controllable T2I models, such as IP-Adapter [44] are in Section B in the Appendix.
@@ 4.4 Analysis and Application
Exploring training settings. We explore several different settings of our method. The details are in Sec. D in the appendix. Some of the takeaways from the experiments are as follows. 1) Model performance is converged around 200k iterations with batch size of 4. 2) Embedding dimension d (i.e., $Z\in\mathbb{R}^{\text{d}}$) needs to be bigger than 128 such as 512 or 2048. 3) The performance difference between isotropic and non-isotropic gaussian prior is marginal.
Controlling multiple attributes. Fig. 4 and Fig. 10 shows the superior performance of Att-Adapter in controlling multi-attributes. More examples and a detailed explanation about the additional attributes yaw and pitch are provided in the appendix (Sec. F.5).
@@
We also show the results of combining LoRA finetuned for a person ID with Att-Adapter in the appendix (Sec. C).
@@ 5 Conclusion and Discussion
Acknowledgement
We would like to express our gratitude to Kalani Murakami for his assistance and support for processing EVOX data during the course of this research.
D.I. acknowledges support from ARL (W911NF-2020-221). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsor.
@@ Appendix A Evaluation Details
Appendix B Additional Related Works
Controllable Text-to-Image Generation
T2I models [31, 30, 27, 34] generate high quality images from text prompts, using large pretrained visual language models like CLIP [29]. However, text only conditioning lacks fine-grained control as text often underspecifies visual details such as object type, perspective and style [17, 45]. To improve control, various approaches incorporate images as additional conditions. ControlNet [46] conditions image generation on inputs like depth, sketches, and semantic maps by training a copy of the diffusion model on these inputs, allowing spatial control. Similarly, T2I-Adapter [26] uses a lightweight adapter for external signals such as images, while preserving the pre-trained model’s capabilities. IP-Adapter [44] adds controllability via cross-attention networks for both text and image inputs, enabling guidance with a reference image.
However, these methods [29, 46, 26] generate images based on a provided example (e.g., body pose) but lack precise control over specific attributes. For instance, with a set of reference faces showing nose widths from narrowest to widest, one might want to generate an image with a specific nose width. Our approach, Att-Adapter, enables precise attribute adjustments using numeric values derived from domain-specific data (e.g., nose width ranges), allowing more accurate modifications within targeted domains.
Appendix C Application results: LoRA
@@
Appendix D Exploring training settings.
@@
Appendix E Additional comparison of Att-Adapter and LoRA
@@
Appendix F Additional Analysis
F.1 Robustness for correlated attributes.
As shown in Table 2 in the main paper, Att-Adapter outperforms baselines in disentangling attributes due to joint attribute learning. To provide further evidence, we performed additional analysis on correlated attributes such as ”mouth width” and ”eye openness”, which typically co-vary through ”smiling” (wider mouth coinciding with narrower eyes). As shown in Fig. 15 left, Att-Adapter can manipulate them independently (wider mouth + and wider eye openness +). More examples showing independent controls over the correlated attributes can be seen in Fig. 15 right.
Figure 15: Disentanglement of correlated attributes by Att-Adapter.
F.2 Dependency on Quality of Attribute Annotations and Paired v.s. Unpaired Data Performance
Att-adapter is specifically designed for scenarios where explicit paired data is unavailable but attribute annotations are present. Such scenarios frequently occur in practice, for example, product images are associated with metadata and manual annotations. In contrast, other methods like ConceptSlider can be effective in scenarios lacking explicit attribute annotations by relying on paired data. Adapting Att-Adapter to such scenarios is nontirivial issue, which can be an interesting future direction.
F.3 Scalability validation
We validated scalability by measuring resource usage with increasing number of attributes. With a batch size of 16 and latent dimension $Z\in\mathbb{R}^{1024}$, using one attribute requires around 23,704MB of VRAM and 97MB of storage. Each additional attribute increases VRAM by only about 338MB ( 1.4% of the initial VRAM) and storage by just 0.004MB ( 0.004% of the initial storage). These results support our claim of handling numerous attributes with negligible increases in memory usage.
F.4 Ablation Study (CVAE v.s. Dropout Regularization)
We explore the effect of dropout in preventing overfitting by conducting additional comparisons. Results in the Table below confirm using dropout (the best rate of 0.25 is reported) indeed reduces identity similarity which indicates improved regularization. However, even at this optimal dropout rate, performance remains worse than our CVAE-based approach. These confirms the effectiveness of CVAE in regularizing Att-Adapter and prevent overfitting.
Naive
w.o. drop with drop Ours
ID sim ($\downarrow$) 28.5% 23.6% 6.2%
F.5 Challenging attributes
We show that Att-Adapter can be used beyond frontal-view images. To show this, we extend our training dataset from the frontal-view face images to the entire FFHQ dataset. We also additionally add three attributes; yaw, pitch, and roll [^6] https://github.com/DCGM/ffhq-features-dataset. The results are shown in Fig. 16. Interestingly, even though Att-Adapter is not specifically designed for understanding 3D domain, we can see that the left-right rotation (i.e., yaw) and the up-down rotation (i.e., pitch) can be controlled. However, we observed that roll is not controllable. To improve this, we believe additional facial dataset with diverse roll information is required as FFHQ face-cropped dataset is face aligned. We also think it would be interesting and powerful if 3D domain knowledge could be aggregated in Att-Adapter which is beyond our research scope.